a mixed solution of potassium hydroxide and sodium carbonate required 15ml of an N/20 hcl solution when titrated with phanolphalein as an indicator .but the same amount of the solution  when titrated with methyl orange as an indicator required 25ml of the same acid .the amount of koh present in th solution.

Grade:11

1 Answers

Vikas TU
4 years ago
Meq of HCl used for mixture using phenolphthalein=15×120
Meq of KOH+12 meq of Na2CO3=34-------(1)
Meq of HCl used for mixture using methyl orange=25×120
(i.e) Meq of KOH+meq of NA2CO3=54------(2)
By equation (1) & (2)
12 Meq of Na2CO3=(54−34)
Meq of Na2CO3=1
w1062×1000=1
w=0.053gNa2CO3
Meq of KOH=54−1=14
w561×1000=14
w=0.014gKOH
Hope this helps